Trace the curve
a2y2=x2(a-x)
1
Expert's answer
2020-04-16T16:24:18-0400

Let "a>0" . The curve "G=\left\{ \left( x,y \right) \ :\ { a }^{ 2 }{ y }^{ 2 }{ =x }^{ 2 }{ (a-x) }^{ } \right\}"  is the union of the function of the graph "y(x)=\frac { x }{ a } \sqrt { a-x }" and "y(x)=-\frac { x }{ a } \sqrt { a-x }" defined on the set "(-\infty ,a]" and symmetric about the "x-" axis. Therefore we draw a curve according to the properties of the function "y(x)=\frac { x }{ a } \sqrt { a-x }"

"1)\quad ay'(x)=\sqrt { a-x } -\frac { x }{ 2\sqrt { a-x } } =\frac { 2a-3x }{ 2\sqrt { a-x } } \ ,x\ <a\quad."

If "x<\frac { 2a }{ 3 }", then "y'(x)>0" and "y(x)\nearrow \quad".

If "\frac { 2a }{ 3 } <x<a" , then "y'(x)<0" and "y(x)\searrow"

2) "2ay"(x)=\ \frac { d }{ dx } \left( \frac { 2a-3x }{ \ \sqrt { a-x } } \right)""=\frac { -3\sqrt { a-x } +\ \frac { (2a-3x) }{ 2\sqrt { a-x } } \ \ \ }{ (a-x) } =" "\frac { -6a+\ 6x+2a-3x }{ 2(a-x)\sqrt { a-x } \ } \ =\frac { -4a+3x }{ 2(a-x)\sqrt { a-x } }"

Since "x<a" , then "-4a+3x<-4a+3a<0" and "y"(x)<0" .So function "y(x)" is convex up.

3) The function graph has no vertical asymptotes , since the function is continuous. There no oblique asymptotes since the "\lim _{ x\rightarrow -\infty \ }{ \frac { f(x) }{ x } } =\lim _{ x\rightarrow -\infty \ }{ \frac { x\sqrt { a-x } \quad }{ ax } = } +\infty"